byeeee:

The population f(x), in millions, of State A of a country after x years is represented by the function shown below: f(x) = 4(1.08)x The graph shows the population g(x), in millions, of State B of the country after x years: Graph of function g of x equals 2 multiplied by 1.08 to the power of x Which conclusion is correct about the population of State A and State B? The original population of State A was double of the original population of State B. The original population of State B was double of the original population of State A. The original population of State A was four times of the original population of State B. The original population of State A was equal to the original population of State B.

Shadow:

\[f(x) = 4(1.08)^{x} \] \[g(x) = 2(1.08)^{x} \]

Shadow:

What do you think the answer is?

byeeee:

a?

Shadow:

Lets do the math.

Shadow:

x represents the time, in years. So 1 would mean the population after 1 year, and so on. If we input 0, we will get the populations with no time having passed. \[f(x) = 4(1.08)^{0}\] Any number raised to the power of 0 is 1. \[f(x) = 4(1) = 4\] \[g(x) = 2(1.08)^{0}\] \[g(x) = 2(1) = 2\]

Shadow:

So what can you confidently say is the answer now?

byeeee:

A!

Shadow:

Correct :)

byeeee:

thank you!

Shadow:

No problem